### v2

# Measurement of the branching fraction of $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ via $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$

Measurement of the branching fraction of $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ via $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$
M. Ablikim1, M. N. Achasov4,c, P. Adlarson75, O. Afedulidis3, X. C. Ai80, R. Aliberti35, A. Amoroso74A,74C, Q. An71,58,a, Y. Bai57, O. Bakina36, I. Balossino29A, Y. Ban46,h, H.-R. Bao63, V. Batozskaya1,44, K. Begzsuren32, N. Berger35, M. Berlowski44, M. Bertani28A, D. Bettoni29A, F. Bianchi74A,74C, E. Bianco74A,74C, A. Bortone74A,74C, I. Boyko36, R. A. Briere5, A. Brueggemann68, H. Cai76, X. Cai1,58, A. Calcaterra28A, G. F. Cao1,63, N. Cao1,63, S. A. Cetin62A, J. F. Chang1,58, G. R. Che43, G. Chelkov36,b, C. Chen43, C. H. Chen9, Chao Chen55, G. Chen1, H. S. Chen1,63, H. Y. Chen20, M. L. Chen1,58,63, S. J. Chen42, S. L. Chen45, S. M. Chen61, T. Chen1,63, X. R. Chen31,63, X. T. Chen1,63, Y. B. Chen1,58, Y. Q. Chen34, Z. J. Chen25,i, Z. Y. Chen1,63, S. K. Choi10A, G. Cibinetto29A, F. Cossio74C, J. J. Cui50, H. L. Dai1,58, J. P. Dai78, A. Dbeyssi18, R.  E. de Boer3, D. Dedovich36, C. Q. Deng72, Z. Y. Deng1, A. Denig35, I. Denysenko36, M. Destefanis74A,74C, F. De Mori74A,74C, B. Ding66,1, X. X. Ding46,h, Y. Ding34, Y. Ding40, J. Dong1,58, L. Y. Dong1,63, M. Y. Dong1,58,63, X. Dong76, M. C. Du1, S. X. Du80, Y. Y. Duan55, Z. H. Duan42, P. Egorov36,b, Y. H. Fan45, J. Fang59, J. Fang1,58, S. S. Fang1,63, W. X. Fang1, Y. Fang1, Y. Q. Fang1,58, R. Farinelli29A, L. Fava74B,74C, F. Feldbauer3, G. Felici28A, C. Q. Feng71,58, J. H. Feng59, Y. T. Feng71,58, M. Fritsch3, C. D. Fu1, J. L. Fu63, Y. W. Fu1,63, H. Gao63, X. B. Gao41, Y. N. Gao46,h, Yang Gao71,58, S. Garbolino74C, I. Garzia29A,29B, L. Ge80, P. T. Ge76, Z. W. Ge42, C. Geng59, E. M. Gersabeck67, A. Gilman69, K. Goetzen13, L. Gong40, W. X. Gong1,58, W. Gradl35, S. Gramigna29A,29B, M. Greco74A,74C, M. H. Gu1,58, Y. T. Gu15, C. Y. Guan1,63, A. Q. Guo31,63, L. B. Guo41, M. J. Guo50, R. P. Guo49, Y. P. Guo12,g, A. Guskov36,b, J. Gutierrez27, K. L. Han63, T. T. Han1, F. Hanisch3, X. Q. Hao19, F. A. Harris65, K. K. He55, K. L. He1,63, F. H. Heinsius3, C. H. Heinz35, Y. K. Heng1,58,63, C. Herold60, T. Holtmann3, P. C. Hong34, G. Y. Hou1,63, X. T. Hou1,63, Y. R. Hou63, Z. L. Hou1, B. Y. Hu59, H. M. Hu1,63, J. F. Hu56,j, S. L. Hu12,g, T. Hu1,58,63, Y. Hu1, G. S. Huang71,58, K. X. Huang59, L. Q. Huang31,63, X. T. Huang50, Y. P. Huang1, Y. S. Huang59, T. Hussain73, F. Hölzken3, N. Hüsken35, N. in der Wiesche68, J. Jackson27, S. Janchiv32, J. H. Jeong10A, Q. Ji1, Q. P. Ji19, W. Ji1,63, X. B. Ji1,63, X. L. Ji1,58, Y. Y. Ji50, X. Q. Jia50, Z. K. Jia71,58, D. Jiang1,63, H. B. Jiang76, P. C. Jiang46,h, S. S. Jiang39, T. J. Jiang16, X. S. Jiang1,58,63, Y. Jiang63, J. B. Jiao50, J. K. Jiao34, Z. Jiao23, S. Jin42, Y. Jin66, M. Q. Jing1,63, X. M. Jing63, T. Johansson75, S. Kabana33, N. Kalantar-Nayestanaki64, X. L. Kang9, X. S. Kang40, M. Kavatsyuk64, B. C. Ke80, V. Khachatryan27, A. Khoukaz68, R. Kiuchi1, O. B. Kolcu62A, B. Kopf3, M. Kuessner3, X. Kui1,63, N.  Kumar26, A. Kupsc44,75, W. Kühn37, J. J. Lane67, L. Lavezzi74A,74C, T. T. Lei71,58, Z. H. Lei71,58, M. Lellmann35, T. Lenz35, C. Li47, C. Li43, C. H. Li39, Cheng Li71,58, D. M. Li80, F. Li1,58, G. Li1, H. B. Li1,63, H. J. Li19, H. N. Li56,j, Hui Li43, J. R. Li61, J. S. Li59, K. Li1, L. J. Li1,63, L. K. Li1, Lei Li48, M. H. Li43, P. R. Li38,k,l, Q. M. Li1,63, Q. X. Li50, R. Li17,31, S. X. Li12, T.  Li50, W. D. Li1,63, W. G. Li1,a, X. Li1,63, X. H. Li71,58, X. L. Li50, X. Y. Li1,63, X. Z. Li59, Y. G. Li46,h, Z. J. Li59, Z. Y. Li78, C. Liang42, H. Liang1,63, H. Liang71,58, Y. F. Liang54, Y. T. Liang31,63, G. R. Liao14, Y. P. Liao1,63, J. Libby26, A.  Limphirat60, C. C. Lin55, D. X. Lin31,63, T. Lin1, B. J. Liu1, B. X. Liu76, C. Liu34, C. X. Liu1, F. Liu1, F. H. Liu53, Feng Liu6, G. M. Liu56,j, H. Liu38,k,l, H. B. Liu15, H. H. Liu1, H. M. Liu1,63, Huihui Liu21, J. B. Liu71,58, J. Y. Liu1,63, K. Liu38,k,l, K. Y. Liu40, Ke Liu22, L. Liu71,58, L. C. Liu43, Lu Liu43, M. H. Liu12,g, P. L. Liu1, Q. Liu63, S. B. Liu71,58, T. Liu12,g, W. K. Liu43, W. M. Liu71,58, X. Liu38,k,l, X. Liu39, Y. Liu80, Y. Liu38,k,l, Y. B. Liu43, Z. A. Liu1,58,63, Z. D. Liu9, Z. Q. Liu50, X. C. Lou1,58,63, F. X. Lu59, H. J. Lu23, J. G. Lu1,58, X. L. Lu1, Y. Lu7, Y. P. Lu1,58, Z. H. Lu1,63, C. L. Luo41, J. R. Luo59, M. X. Luo79, T. Luo12,g, X. L. Luo1,58, X. R. Lyu63, Y. F. Lyu43, F. C. Ma40, H. Ma78, H. L. Ma1, J. L. Ma1,63, L. L. Ma50, M. M. Ma1,63, Q. M. Ma1, R. Q. Ma1,63, T. Ma71,58, X. T. Ma1,63, X. Y. Ma1,58, Y. Ma46,h, Y. M. Ma31, F. E. Maas18, M. Maggiora74A,74C, S. Malde69, Y. J. Mao46,h, Z. P. Mao1, S. Marcello74A,74C, Z. X. Meng66, J. G. Messchendorp13,64, G. Mezzadri29A, H. Miao1,63, T. J. Min42, R. E. Mitchell27, X. H. Mo1,58,63, B. Moses27, N. Yu. Muchnoi4,c, J. Muskalla35, Y. Nefedov36, F. Nerling18,e, L. S. Nie20, I. B. Nikolaev4,c, Z. Ning1,58, S. Nisar11,m, Q. L. Niu38,k,l, W. D. Niu55, Y. Niu 50, S. L. Olsen63, Q. Ouyang1,58,63, S. Pacetti28B,28C, X. Pan55, Y. Pan57, A.  Pathak34, Y. P. Pei71,58, M. Pelizaeus3, H. P. Peng71,58, Y. Y. Peng38,k,l, K. Peters13,e, J. L. Ping41, R. G. Ping1,63, S. Plura35, V. Prasad33, F. Z. Qi1, H. Qi71,58, H. R. Qi61, M. Qi42, T. Y. Qi12,g, S. Qian1,58, W. B. Qian63, C. F. Qiao63, X. K. Qiao80, J. J. Qin72, L. Q. Qin14, L. Y. Qin71,58, X. P. Qin12,g, X. S. Qin50, Z. H. Qin1,58, J. F. Qiu1, Z. H. Qu72, C. F. Redmer35, K. J. Ren39, A. Rivetti74C, M. Rolo74C, G. Rong1,63, Ch. Rosner18, S. N. Ruan43, N. Salone44, A. Sarantsev36,d, Y. Schelhaas35, K. Schoenning75, M. Scodeggio29A, K. Y. Shan12,g, W. Shan24, X. Y. Shan71,58, Z. J. Shang38,k,l, J. F. Shangguan16, L. G. Shao1,63, M. Shao71,58, C. P. Shen12,g, H. F. Shen1,8, W. H. Shen63, X. Y. Shen1,63, B. A. Shi63, H. Shi71,58, H. C. Shi71,58, J. L. Shi12,g, J. Y. Shi1, Q. Q. Shi55, S. Y. Shi72, X. Shi1,58, J. J. Song19, T. Z. Song59, W. M. Song34,1, Y.  J. Song12,g, Y. X. Song46,h,n, S. Sosio74A,74C, S. Spataro74A,74C, F. Stieler35, Y. J. Su63, G. B. Sun76, G. X. Sun1, H. Sun63, H. K. Sun1, J. F. Sun19, K. Sun61, L. Sun76, S. S. Sun1,63, T. Sun51,f, W. Y. Sun34, Y. Sun9, Y. J. Sun71,58, Y. Z. Sun1, Z. Q. Sun1,63, Z. T. Sun50, C. J. Tang54, G. Y. Tang1, J. Tang59, M. Tang71,58, Y. A. Tang76, L. Y. Tao72, Q. T. Tao25,i, M. Tat69, J. X. Teng71,58, V. Thoren75, W. H. Tian59, Y. Tian31,63, Z. F. Tian76, I. Uman62B, Y. Wan55, S. J. Wang 50, B. Wang1, B. L. Wang63, Bo Wang71,58, D. Y. Wang46,h, F. Wang72, H. J. Wang38,k,l, J. J. Wang76, J. P. Wang 50, K. Wang1,58, L. L. Wang1, M. Wang50, N. Y. Wang63, S. Wang12,g, S. Wang38,k,l, T.  Wang12,g, T. J. Wang43, W. Wang59, W.  Wang72, W. P. Wang35,71,o, X. Wang46,h, X. F. Wang38,k,l, X. J. Wang39, X. L. Wang12,g, X. N. Wang1, Y. Wang61, Y. D. Wang45, Y. F. Wang1,58,63, Y. L. Wang19, Y. N. Wang45, Y. Q. Wang1, Yaqian Wang17, Yi Wang61, Z. Wang1,58, Z. L.  Wang72, Z. Y. Wang1,63, Ziyi Wang63, D. H. Wei14, F. Weidner68, S. P. Wen1, Y. R. Wen39, U. Wiedner3, G. Wilkinson69, M. Wolke75, L. Wollenberg3, C. Wu39, J. F. Wu1,8, L. H. Wu1, L. J. Wu1,63, X. Wu12,g, X. H. Wu34, Y. Wu71,58, Y. H. Wu55, Y. J. Wu31, Z. Wu1,58, L. Xia71,58, X. M. Xian39, B. H. Xiang1,63, T. Xiang46,h, D. Xiao38,k,l, G. Y. Xiao42, S. Y. Xiao1, Y.  L. Xiao12,g, Z. J. Xiao41, C. Xie42, X. H. Xie46,h, Y. Xie50, Y. G. Xie1,58, Y. H. Xie6, Z. P. Xie71,58, T. Y. Xing1,63, C. F. Xu1,63, C. J. Xu59, G. F. Xu1, H. Y. Xu66,2,p, M. Xu71,58, Q. J. Xu16, Q. N. Xu30, W. Xu1, W. L. Xu66, X. P. Xu55, Y. C. Xu77, Z. S. Xu63, F. Yan12,g, L. Yan12,g, W. B. Yan71,58, W. C. Yan80, X. Q. Yan1, H. J. Yang51,f, H. L. Yang34, H. X. Yang1, T. Yang1, Y. Yang12,g, Y. F. Yang1,63, Y. F. Yang43, Y. X. Yang1,63, Z. W. Yang38,k,l, Z. P. Yao50, M. Ye1,58, M. H. Ye8, J. H. Yin1, Z. Y. You59, B. X. Yu1,58,63, C. X. Yu43, G. Yu1,63, J. S. Yu25,i, T. Yu72, X. D. Yu46,h, Y. C. Yu80, C. Z. Yuan1,63, J. Yuan34, J. Yuan45, L. Yuan2, S. C. Yuan1,63, Y. Yuan1,63, Z. Y. Yuan59, C. X. Yue39, A. A. Zafar73, F. R. Zeng50, S. H.  Zeng72, X. Zeng12,g, Y. Zeng25,i, Y. J. Zeng59, Y. J. Zeng1,63, X. Y. Zhai34, Y. C. Zhai50, Y. H. Zhan59, A. Q. Zhang1,63, B. L. Zhang1,63, B. X. Zhang1, D. H. Zhang43, G. Y. Zhang19, H. Zhang80, H. Zhang71,58, H. C. Zhang1,58,63, H. H. Zhang34, H. H. Zhang59, H. Q. Zhang1,58,63, H. R. Zhang71,58, H. Y. Zhang1,58, J. Zhang80, J. Zhang59, J. J. Zhang52, J. L. Zhang20, J. Q. Zhang41, J. S. Zhang12,g, J. W. Zhang1,58,63, J. X. Zhang38,k,l, J. Y. Zhang1, J. Z. Zhang1,63, Jianyu Zhang63, L. M. Zhang61, Lei Zhang42, P. Zhang1,63, Q. Y. Zhang34, R. Y. Zhang38,k,l, S. H. Zhang1,63, Shulei Zhang25,i, X. D. Zhang45, X. M. Zhang1, X. Y. Zhang50, Y.  Zhang72, Y. Zhang1, Y.  T. Zhang80, Y. H. Zhang1,58, Y. M. Zhang39, Yan Zhang71,58, Z. D. Zhang1, Z. H. Zhang1, Z. L. Zhang34, Z. Y. Zhang76, Z. Y. Zhang43, Z. Z.  Zhang45, G. Zhao1, J. Y. Zhao1,63, J. Z. Zhao1,58, L. Zhao1, Lei Zhao71,58, M. G. Zhao43, N. Zhao78, R. P. Zhao63, S. J. Zhao80, Y. B. Zhao1,58, Y. X. Zhao31,63, Z. G. Zhao71,58, A. Zhemchugov36,b, B. Zheng72, B. M. Zheng34, J. P. Zheng1,58, W. J. Zheng1,63, Y. H. Zheng63, B. Zhong41, X. Zhong59, H.  Zhou50, J. Y. Zhou34, L. P. Zhou1,63, S.  Zhou6, X. Zhou76, X. K. Zhou6, X. R. Zhou71,58, X. Y. Zhou39, Y. Z. Zhou12,g, J. Zhu43, K. Zhu1, K. J. Zhu1,58,63, K. S. Zhu12,g, L. Zhu34, L. X. Zhu63, S. H. Zhu70, T. J. Zhu12,g, W. D. Zhu41, Y. C. Zhu71,58, Z. A. Zhu1,63, J. H. Zou1, J. Zu71,58
(BESIII Collaboration)
1 Institute of High Energy Physics, Beijing 100049, People’s Republic of China
2 Beihang University, Beijing 100191, People’s Republic of China
3 Bochum Ruhr-University, D-44780 Bochum, Germany
4 Budker Institute of Nuclear Physics SB RAS (BINP), Novosibirsk 630090, Russia
5 Carnegie Mellon University, Pittsburgh, Pennsylvania 15213, USA
6 Central China Normal University, Wuhan 430079, People’s Republic of China
7 Central South University, Changsha 410083, People’s Republic of China
8 China Center of Advanced Science and Technology, Beijing 100190, People’s Republic of China
9 China University of Geosciences, Wuhan 430074, People’s Republic of China
10 Chung-Ang University, Seoul, 06974, Republic of Korea
11 COMSATS University Islamabad, Lahore Campus, Defence Road, Off Raiwind Road, 54000 Lahore, Pakistan
12 Fudan University, Shanghai 200433, People’s Republic of China
13 GSI Helmholtzcentre for Heavy Ion Research GmbH, D-64291 Darmstadt, Germany
14 Guangxi Normal University, Guilin 541004, People’s Republic of China
15 Guangxi University, Nanning 530004, People’s Republic of China
16 Hangzhou Normal University, Hangzhou 310036, People’s Republic of China
17 Hebei University, Baoding 071002, People’s Republic of China
18 Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
19 Henan Normal University, Xinxiang 453007, People’s Republic of China
20 Henan University, Kaifeng 475004, People’s Republic of China
21 Henan University of Science and Technology, Luoyang 471003, People’s Republic of China
22 Henan University of Technology, Zhengzhou 450001, People’s Republic of China
23 Huangshan College, Huangshan 245000, People’s Republic of China
24 Hunan Normal University, Changsha 410081, People’s Republic of China
25 Hunan University, Changsha 410082, People’s Republic of China
26 Indian Institute of Technology Madras, Chennai 600036, India
27 Indiana University, Bloomington, Indiana 47405, USA
28 INFN Laboratori Nazionali di Frascati , (A)INFN Laboratori Nazionali di Frascati, I-00044, Frascati, Italy; (B)INFN Sezione di Perugia, I-06100, Perugia, Italy; (C)University of Perugia, I-06100, Perugia, Italy
29 INFN Sezione di Ferrara, (A)INFN Sezione di Ferrara, I-44122, Ferrara, Italy; (B)University of Ferrara, I-44122, Ferrara, Italy
30 Inner Mongolia University, Hohhot 010021, People’s Republic of China
31 Institute of Modern Physics, Lanzhou 730000, People’s Republic of China
32 Institute of Physics and Technology, Peace Avenue 54B, Ulaanbaatar 13330, Mongolia
33 Instituto de Alta Investigación, Universidad de Tarapacá, Casilla 7D, Arica 1000000, Chile
34 Jilin University, Changchun 130012, People’s Republic of China
35 Johannes Gutenberg University of Mainz, Johann-Joachim-Becher-Weg 45, D-55099 Mainz, Germany
36 Joint Institute for Nuclear Research, 141980 Dubna, Moscow region, Russia
37 Justus-Liebig-Universitaet Giessen, II. Physikalisches Institut, Heinrich-Buff-Ring 16, D-35392 Giessen, Germany
38 Lanzhou University, Lanzhou 730000, People’s Republic of China
39 Liaoning Normal University, Dalian 116029, People’s Republic of China
40 Liaoning University, Shenyang 110036, People’s Republic of China
41 Nanjing Normal University, Nanjing 210023, People’s Republic of China
42 Nanjing University, Nanjing 210093, People’s Republic of China
43 Nankai University, Tianjin 300071, People’s Republic of China
44 National Centre for Nuclear Research, Warsaw 02-093, Poland
45 North China Electric Power University, Beijing 102206, People’s Republic of China
46 Peking University, Beijing 100871, People’s Republic of China
47 Qufu Normal University, Qufu 273165, People’s Republic of China
48 Renmin University of China, Beijing 100872, People’s Republic of China
49 Shandong Normal University, Jinan 250014, People’s Republic of China
50 Shandong University, Jinan 250100, People’s Republic of China
51 Shanghai Jiao Tong University, Shanghai 200240, People’s Republic of China
52 Shanxi Normal University, Linfen 041004, People’s Republic of China
53 Shanxi University, Taiyuan 030006, People’s Republic of China
54 Sichuan University, Chengdu 610064, People’s Republic of China
55 Soochow University, Suzhou 215006, People’s Republic of China
56 South China Normal University, Guangzhou 510006, People’s Republic of China
57 Southeast University, Nanjing 211100, People’s Republic of China
58 State Key Laboratory of Particle Detection and Electronics, Beijing 100049, Hefei 230026, People’s Republic of China
59 Sun Yat-Sen University, Guangzhou 510275, People’s Republic of China
60 Suranaree University of Technology, University Avenue 111, Nakhon Ratchasima 30000, Thailand
61 Tsinghua University, Beijing 100084, People’s Republic of China
62 Turkish Accelerator Center Particle Factory Group, (A)Istinye University, 34010, Istanbul, Turkey; (B)Near East University, Nicosia, North Cyprus, 99138, Mersin 10, Turkey
63 University of Chinese Academy of Sciences, Beijing 100049, People’s Republic of China
64 University of Groningen, NL-9747 AA Groningen, The Netherlands
65 University of Hawaii, Honolulu, Hawaii 96822, USA
66 University of Jinan, Jinan 250022, People’s Republic of China
67 University of Manchester, Oxford Road, Manchester, M13 9PL, United Kingdom
68 University of Muenster, Wilhelm-Klemm-Strasse 9, 48149 Muenster, Germany
69 University of Oxford, Keble Road, Oxford OX13RH, United Kingdom
70 University of Science and Technology Liaoning, Anshan 114051, People’s Republic of China
71 University of Science and Technology of China, Hefei 230026, People’s Republic of China
72 University of South China, Hengyang 421001, People’s Republic of China
73 University of the Punjab, Lahore-54590, Pakistan
74 University of Turin and INFN, (A)University of Turin, I-10125, Turin, Italy; (B)University of Eastern Piedmont, I-15121, Alessandria, Italy; (C)INFN, I-10125, Turin, Italy
75 Uppsala University, Box 516, SE-75120 Uppsala, Sweden
76 Wuhan University, Wuhan 430072, People’s Republic of China
77 Yantai University, Yantai 264005, People’s Republic of China
78 Yunnan University, Kunming 650500, People’s Republic of China
79 Zhejiang University, Hangzhou 310027, People’s Republic of China
80 Zhengzhou University, Zhengzhou 450001, People’s Republic of China
a Deceased
b Also at the Moscow Institute of Physics and Technology, Moscow 141700, Russia
c Also at the Novosibirsk State University, Novosibirsk, 630090, Russia
d Also at the NRC ”Kurchatov Institute”, PNPI, 188300, Gatchina, Russia
e Also at Goethe University Frankfurt, 60323 Frankfurt am Main, Germany
f Also at Key Laboratory for Particle Physics, Astrophysics and Cosmology, Ministry of Education; Shanghai Key Laboratory for Particle Physics and Cosmology; Institute of Nuclear and Particle Physics, Shanghai 200240, People’s Republic of China
g Also at Key Laboratory of Nuclear Physics and Ion-beam Application (MOE) and Institute of Modern Physics, Fudan University, Shanghai 200443, People’s Republic of China
h Also at State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing 100871, People’s Republic of China
i Also at School of Physics and Electronics, Hunan University, Changsha 410082, China
j Also at Guangdong Provincial Key Laboratory of Nuclear Science, Institute of Quantum Matter, South China Normal University, Guangzhou 510006, China
k Also at MOE Frontiers Science Center for Rare Isotopes, Lanzhou University, Lanzhou 730000, People’s Republic of China
l Also at Lanzhou Center for Theoretical Physics, Lanzhou University, Lanzhou 730000, People’s Republic of China
m Also at the Department of Mathematical Sciences, IBA, Karachi 75270, Pakistan
n Also at Ecole Polytechnique Federale de Lausanne (EPFL), CH-1015 Lausanne, Switzerland
o Also at Helmholtz Institute Mainz, Staudinger Weg 18, D-55099 Mainz, Germany
p Also at School of Physics, Beihang University, Beijing 100191 , China
Abstract
Based on $10.64~{}\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at center-of-mass energies between 4.237 and 4.699 GeV with the BESIII detector, we study the leptonic $D^{+}_{s}$ decays using the $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ process. The branching fractions of $D_{s}^{+}\to\ell^{+}\nu_{\ell}\,(\ell=\mu,\tau)$ are measured to be $\mathcal{B}(D_{s}^{+}\to\mu^{+}\nu_{\mu})=(0.547\pm 0.026_{\rm stat}\pm 0.016_{\rm syst})\%$ and $\mathcal{B}(D_{s}^{+}\to\tau^{+}\nu_{\tau})=(5.60\pm 0.16_{\rm stat}\pm 0.20_{\rm syst})\%$, respectively. The product of the decay constant and Cabibbo-Kobayashi-Maskawa matrix element $|V_{cs}|$ is determined to be $f_{D_{s}^{+}}|V_{cs}|=(246.5\pm 5.9_{\rm stat}\pm 3.6_{\rm syst}\pm 0.5_{\rm input})_{\mu\nu}~{}\mathrm{MeV}$ and $f_{D_{s}^{+}}|V_{cs}|=(252.7\pm 3.6_{\rm stat}\pm 4.5_{\rm syst}\pm 0.6_{\rm input}))_{\tau\nu}~{}\mathrm{MeV}$, respectively. Taking the value of $|V_{cs}|$ from a global fit in the Standard Model, we obtain ${f_{D^{+}_{s}}}=(252.8\pm 6.0_{\rm stat}\pm 3.7_{\rm syst}\pm 0.6_{\rm input})_{\mu\nu}$ MeV and ${f_{D^{+}_{s}}}=(259.2\pm 3.6_{\rm stat}\pm 4.5_{\rm syst}\pm 0.6_{\rm input})_{\tau\nu}$ MeV, respectively. Conversely, taking the value for $f_{D_{s}^{+}}$ from the latest lattice quantum chromodynamics calculation, we obtain $|V_{cs}|=(0.986\pm 0.023_{\rm stat}\pm 0.014_{\rm syst}\pm 0.003_{\rm input})_{\mu\nu}$ and $|V_{cs}|=(1.011\pm 0.014_{\rm stat}\pm 0.018_{\rm syst}\pm 0.003_{\rm input})_{\tau\nu}$, respectively.
I Introduction
Experimental studies of the decays of the $D^{+}_{s}$ are important to understand weak and strong interactions in the charm decays. In the Standard Model (SM), the weak and strong effects in leptonic $D_{s}^{+}$ decays can be well-separated. The partial width of the decay $D_{s}^{+}\to\ell^{+}\nu_{\ell}\,(\ell=\mu,\tau)$ is given by decayrate
$\Gamma_{D_{s}^{+}\to\ell^{+}\nu_{\ell}}=\frac{G^{2}_{F}}{8\pi}f^{2}_{D_{s}^{+}}|V_{cs}|^{2}m^{2}_{\ell}m_{D_{s}^{+}}\left(1-\frac{m^{2}_{\ell}}{m^{2}_{D_{s}^{+}}}\right)^{2},$ (1)
where $G_{F}$ is the Fermi coupling constant, $f_{D_{s}^{+}}$ is the $D_{s}^{+}$ decay constant, $|V_{cs}|$ is the magnitude of the $c\to s$ Cabibbo-Kobayashi-Maskawa (CKM) matrix element pdg2024 , $m_{\ell}$ is the lepton mass, and $m_{D_{s}^{+}}$ is the $D_{s}^{+}$ meson mass. Using the measured branching fraction (BF) of $D_{s}^{+}\to\ell^{+}\nu_{\ell}$, the product of $f_{D_{s}^{+}}|V_{cs}|$ can be determined. By taking the latest $f_{D_{s}^{+}}$ prd98_074512 calculated by lattice quantum chromodynamics (LQCD), one can determine $|V_{cs}|$, which is an essential input for testing CKM matrix unitarity. Conversely, taking $|V_{cs}|$ from the SM global fit, one can extract $f_{D_{s}^{+}}$, which is a crucial check of LQCD calculations prd98_074512 ; fds1 ; fds2 ; fds3 ; fds4 ; fds5 . In addition, the BF ratio of $D_{s}^{+}\to\tau^{+}\nu_{\tau}$ and $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ provides an important test of $\tau-\mu$ lepton-flavor universality.
In recent years, many studies of $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ have been performed by the CLEO cleo2009 ; cleo2009a ; cleo2009b , BaBar babar2010 , Belle belle2013 , and BESIII bes4009 ; bes3_Ds_muv ; bes3_Ds_muv1 ; bes3_Ds_tauv1 ; bes3_Ds_tauv2 ; bes3_Ds_tauv3 ; xiechenpiv ; llcmvv ; Ke:2023qzc ; Li:2021iwf experiments. The BESIII Collaboration has reported experimental studies of $D_{s}^{+}\to\ell^{+}\nu_{\ell}$ using the $e^{+}e^{-}\to D^{+}_{s}D^{-}_{s}$ and $e^{+}e^{-}\to D^{\pm}_{s}D^{*\mp}_{s}$ processes. These studies are based on 0.48 fb-1 and 7.33 fb-1 of $e^{+}e^{-}$ collision data taken at the center-of-mass energies($E_{\rm cm}$) of $\sqrt{s}=$ 4.009 GeV bes4009 and $4.128-4.226$ GeV bes3_Ds_muv ; bes3_Ds_muv1 ; bes3_Ds_tauv1 ; bes3_Ds_tauv2 ; bes3_Ds_tauv3 ; xiechenpiv ; llcmvv , respectively. The latter ones are the most precise measurements to date. In this paper, we perform new measurements of the BFs of $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ via the $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ process. This analysis utilizes 10.64 fb-1 of $e^{+}e^{-}$ collision data collected at center-of-mass energies between $\sqrt{s}=4.237$ and $4.699$ GeV. Notably, this is the first time the $e^{+}e^{-}\to D_{s}^{*+}D_{s}^{*-}$ process has been used to measure $D_{s}^{+}$ leptonic decays. Throughout this paper, charge-conjugation is always implied and $\rho$ denotes the $\rho(770)$ meson.
II BESIII detector and MC simulation
[…]
Simulated samples produced with the geant4-based geant4 MC package, which includes the geometric description of the BESIII detector and the detector response, are used to determine the detection efficiency and to estimate the backgrounds. The simulation includes the beam-energy spread and initial-state radiation in the $e^{+}e^{-}$ annihilations modeled with the generator kkmc kkmc . An inclusive MC sample with a luminosity of 20 times that of the data is produced at center-of-mass energies between 4.237 and 4.699 GeV. It includes open-charm processes, initial state radiation production of $\psi(3770)$, $\psi(3686)$ and $J/\psi$, $q\bar{q}$ $(q=u,d,s)$ continuum processes, along with Bhabha scattering, $\mu^{+}\mu^{-}$, $\tau^{+}\tau^{-}$, and $\gamma\gamma$ events. In the simulation, the production of open-charm processes directly via $e^{+}e^{-}$ annihilations are modeled with the generator conexc conexc . The known decay modes are modeled with evtgen evtgen using the BFs taken from the Particle Data Group (PDG) pdg2024 , and the remaining unknown decays of the charmonium states are modeled by lundcharm lundcharm . Final-state radiation is incorporated using the photos package photos . The input energy-dependent Born cross section for $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ is based on the BESIII measurement crsDssDss .
III Analysis method
In the $e^{+}e^{-}\to D^{*+}_{s}D^{*-}_{s}$ process, the $D_{s}$ mesons are produced via $D_{s}^{*}\to\gamma(\pi^{0})D_{s}$. We fully reconstruct the transition $\gamma(\pi^{0})$ and the $D_{s}^{-}$ meson in one of several hadronic decay modes; successful cases are referred to as single-tag candidates. When the single-tag $D_{s}^{*-}$ and the signal $D_{s}^{*+}$ decay of interest are simultaneously reconstructed, we obtain the so-called double-tag candidates. The BF of the signal decay is determined by
$\mathcal{B}_{\rm sig}=\frac{N_{\rm DT}}{N_{\rm ST}\cdot\bar{\epsilon}_{{\rm sig}}}.$ (2)
Here, $N_{\rm DT}$ is the double-tag yield in data; $N_{\rm ST}=\Sigma_{i,j}N_{\rm ST}^{i,j}$ is the total single-tag yield in data summing over the tag mode $i$ and the energy point $j$; the $\bar{\epsilon}_{\rm sig}$ is the averaged efficiency of the signal decay, and estimated by $\bar{\epsilon}_{\rm sig}=\sum_{j}\left[\sum_{i}\left(\frac{N_{\rm ST}^{i}}{N_{\rm ST}^{j}}\cdot\frac{\epsilon_{\rm DT}^{i}}{\epsilon_{\rm ST}^{i}}\right)\cdot\frac{N_{\rm ST}^{j}}{N_{\rm ST}}\right]$, where $\epsilon_{\rm DT}^{i}$ and $\epsilon_{\rm ST}^{i}$ are the detection efficiencies of the double-tag and single-tag candidates for the $i$-th tag mode, respectively. The efficiencies include the BFs of the daughter particle decays.
IV Single-tag $D^{*-}_{s}$ candidates
The single-tag $D^{*-}_{s}$ candidates are formed from the selected $D^{-}_{s}$ candidates and a transition $\gamma(\pi^{0})$. The $D^{-}_{s}$ candidates are reconstructed from the fourteen hadronic decay modes, including $D^{-}_{s}\to K^{+}K^{-}\pi^{-}$, $K^{+}K^{-}\pi^{-}\pi^{0}$, $\pi^{+}\pi^{-}\pi^{-}$, $K^{0}_{S}K^{-}$, $K^{0}_{S}K^{-}\pi^{0}$, $K^{-}\pi^{+}\pi^{+}$, $K^{0}_{S}K^{0}_{S}\pi^{-}$, $K^{0}_{S}K^{+}\pi^{-}\pi^{-}$, $K^{0}_{S}K^{-}\pi^{+}\pi^{-}$, $\eta_{\gamma\gamma}\pi^{-}$, $\eta_{\pi^{0}\pi^{+}\pi^{-}}\pi^{-}$, $\eta^{\prime}_{\pi^{+}\pi^{-}\eta}\pi^{-}$, $\eta^{\prime}_{\gamma\rho^{0}}\pi^{-}$, and $\eta_{\gamma\gamma}\rho^{-}$. Here, the subscripts of the $\eta$ and $\eta^{\prime}$ represent the decay modes used to reconstruct those states.
[…]
Each $K_{S}^{0}$ candidate is reconstructed from two oppositely charged tracks satisfying $|V_{z}|<$ 20 cm. The two charged tracks are assigned as $\pi^{+}\pi^{-}$ without imposing further PID criteria. They are constrained to originate from a common vertex and are required to have an invariant mass within $|M_{\pi^{+}\pi^{-}}-m_{K_{S}^{0}}|<$ 12 MeV$/c^{2}$, where $m_{K_{S}^{0}}$ is the $K^{0}_{S}$ nominal mass pdg2024 . The decay length of the $K^{0}_{S}$ candidate is required to be greater than twice the vertex resolution away from the IP.
Photon candidates are identified using isolated showers in the EMC. The deposited energy of each shower must be more than 25 MeV in the barrel region ($|\cos\theta|<0.80$) and more than 50 MeV in the end cap region ($0.920.86<|\cos\theta|<$ 0.92). To exclude showers that originate from charged tracks, the angle subtended by the EMC shower and the position of the closest charged track at the EMC must be greater than 10 degrees as measured from the IP. To suppress electronic noise and showers unrelated to the event, the difference between the EMC time and the event start time is required to be within [0, 700] ns.
The $\pi^{0}$ and $\eta$ mesons are reconstructed from photon pairs. To form $\pi^{0}$ and $\eta$ candidates, the invariant masses of the selected photon pairs are required to be within the $M_{\gamma\gamma}$ intervals (0.115, 0.150)0.1150.150(0.115,\,0.150)( 0.115 , 0.150 ) GeV$/c^{2}$ and (0.500, 0.570)0.5000.570(0.500,\,0.570)( 0.500 , 0.570 ) GeV$/c^{2}$, respectively. To improve momentum resolution and suppress background, a kinematic fit is imposed on each photon pair to constrain their invariant mass to the nominal $\pi^{0}$ or $\eta$ mass pdg2024 . The $\chi^{2}$ of this kinematic fit is required to be less than 20.
The $\eta$ candidates are also formed from $\pi^{0}\pi^{+}\pi^{-}$ combinations with invariant masses in the interval (0.530, $(0.530,\,0.570)~{}\mathrm{GeV}/c^{2}$. The $\eta^{\prime}$ candidates are formed from the $\eta\pi^{+}\pi^{-}$ and $\gamma\rho^{0}$ combinations with invariant masses in the intervals (0.946, 0.970)0.9460.970(0.946,\,0.970)( 0.946 , 0.970 ) GeV/$c^{2}$ and (0.940, $(0.940,\,0.976)~{}\mathrm{GeV}/c^{2}$, respectively. In addition, the minimum energy of the $\gamma$ from $\eta^{\prime}\to\gamma\rho^{0}$ decays must be greater than 0.1 GeV. The $\rho^{0}$ and $\rho^{+}$ candidates are reconstructed from the $\pi^{+}\pi^{-}$ and $\pi^{+}\pi^{0}$ combinations with invariant masses within the common interval (0.570, $(0.570,\,0.970)~{}\mathrm{GeV}/c^{2}$.
To reject the peaking background events from $D_{s}^{-}\to K^{0}_{S}(\to\pi^{+}\pi^{-})\pi^{-}$ and $D_{s}^{-}\to K^{0}_{S}(\to\pi^{+}\pi^{-})K^{-}$ in the tag modes of $D_{s}^{-}\to\pi^{+}\pi^{-}\pi^{-}$ and $D_{s}^{-}\to K^{+}\pi^{+}\pi^{-}$, we require that the invariant mass of any $\pi^{+}\pi^{-}$ combination satisfy $|M_{\pi^{+}\pi^{-}}-m_{K_{S}^{0}}|>0.03$ GeV/$c^{2}$.
Figure 1: Fits to the $M_{\rm BC}$ distributions of the single-tag $D^{*-}_{s}$ candidates, where the points with error bars are data at 4.260 GeV, the blue solid curves show the best fits, and the red dashed curves show the fitted combinatorial background shapes. shapes. The pairs of arrows denote the $M_{\rm BC}$ signal window.
The tagged $D^{-}_{s}$ candidates are required to fall in the $M_{D_{s}}$ mass windows shown in the second column of Table 1, following Refs. panx1 ; panx2 . To further distinguish the single-tag $D^{*-}_{s}$ from combinatorial background, we use two kinematic variables: the energy difference defined as
$\Delta E=E_{\rm tag}-E_{\rm beam},$ (3)
and the beam-constrained mass defined as
$M_{\rm BC}=\sqrt{E^{2}_{\rm beam}/c^{4}-|\vec{p}_{\rm tag}|^{2}/c^{2}},$ (4)
where $E_{\rm beam}$ is the beam energy, and $E_{\rm tag}$ and $\vec{p}_{\rm tag}$ are the energy and momentum of the single-tag $D^{*-}_{s}$ candidate in the rest frame of the initial $e^{+}e^{-}$ beams.
The $\Delta E$ values for each tag must be within the ranges listed in Table 1. If multiple $\gamma/\pi^{0}$ or $D_{s}$ combinations remain, we keep only the candidates with the minimum $|\Delta E|$ for each tag mode and each $D_{s}^{*}$ charge from a given event. For each tag mode, the single-tag yield is obtained by a fit to the corresponding $M_{\rm BC}$ spectrum. The signal is described by the simulated shape convolved with a Gaussian function representing the difference in resolution between data and simulation. Because the endpoint above 4.450 GeV is far away from the nominal $D^{*}_{s}$ mass, the data at all energy points above 4.450 GeV are combined into one data set. For the data sets taken below and above 4.450 GeV, the non-peaking background shapes are modeled by an ARGUS function argus and a second order Chebychev polynomial function, respectively, which have been validated by analyzing the inclusive MC sample. As an example, the $M_{\rm BC}$ distributions of accepted single-tag candidates for various tag modes and the fit results at 4.260 GeV are shown in Fig. 1. The candidates in the signal regions, denoted as the red arrows in each sub-figure, are kept for further analysis. The resulting single-tag yields ($N_{\rm ST}^{i}$) for various tag modes and the corresponding single-tag efficiencies ($\epsilon_{\rm ST}^{i}$) at 4.260 GeV are shown in Table 1. Information for data from other energies is given in Appendix A. The total single-tag yields at the different energy points are summarized in Table 2.
Table 1: The $M_{D^{-}_{s}}$ requirements, $\Delta E$ requirements, single-tag yields ($N_{\rm ST}$), single-tag efficiencies ($\epsilon_{\rm ST}$), and various double-tag efficiencies for each tag mode at 4.260 GeV. The $\epsilon^{\mu_{a}\nu}_{\rm DT}$ and $\epsilon^{\mu_{b}\nu}_{\rm DT}$ correspond to the double-tag efficiencies for $D^{+}_{s}\to\mu^{+}_{a}\nu_{\mu}$ and $D^{+}_{s}\to\mu^{+}_{b}\nu_{\mu}$, while $\epsilon^{\tau_{e}\nu}_{\rm DT}$, $\epsilon^{\tau_{\mu}\nu}_{\rm DT}$, $\epsilon^{\tau_{\pi}\nu}_{\rm DT}$, and $\epsilon^{\tau_{\rho}\nu}_{\rm DT}$ correspond to the double-tag efficiencies for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, respectively. The Muon candidates identified with (without) MUC information are denoted by $\mu_{a}$ ($\mu_{b}$). The efficiencies include the BFs of $D_{s}^{*+}$ and $\tau^{+}$ decays. The uncertainties are statistical only.
$D^{-}_{s}$ tag mode $M_{D^{-}_{s}}$ (GeV/$c^{2}$) $\Delta E$ (MeV) $N_{\rm ST}$ $\epsilon_{\rm ST}$ (%) $\epsilon^{\mu_{a}\nu}_{\rm DT}$ (%) $\epsilon^{\mu_{b}\nu}_{\rm DT}$ (%) $\epsilon^{\tau_{e}\nu}_{\rm DT}$ (%) $\epsilon^{\tau_{\mu}\nu}_{\rm DT}$ (%) $\epsilon^{\tau_{\pi}\nu}_{\rm DT}$ (%) $\epsilon^{\tau_{\rho}\nu}_{\rm DT}$ (%)
$K^{+}K^{-}\pi^{-}$ $(1.950,1.986)$ $(-31,26)$ $7454$$\pm 125$ $19.67$$\pm 0.07$ $14.43$$\pm 0.07$ $16.89$$\pm 0.09$ $3.54$$\pm 0.02$ $1.53$$\pm 0.01$ $1.62$$\pm 0.01$ $1.07$$\pm 0.01$
$K^{+}K^{-}\pi^{-}\pi^{0}$ $(1.947,1.982)$ $(-38,29)$ $2186$$\pm 108$ $5.15$$\pm 0.05$ $5.08$$\pm 0.06$ $5.53$$\pm 0.06$ $1.21$$\pm 0.01$ $0.55$$\pm 0.01$ $0.52$$\pm 0.01$ $0.28$$\pm 0.01$
$\pi^{-}\pi^{+}\pi^{-}$ $(1.952,1.984)$ $(-34,28)$ $1929$$\pm 99$ $25.73$$\pm 0.26$ $18.18$$\pm 0.08$ $24.60$$\pm 0.10$ $4.32$$\pm 0.02$ $1.84$$\pm 0.01$ $2.35$$\pm 0.01$ $1.76$$\pm 0.02$
$K_{S}^{0}K^{-}$ $(1.948,1.991)$ $(-33,30)$ $1649$$\pm 53$ $22.97$$\pm 0.16$ $16.44$$\pm 0.09$ $21.80$$\pm 0.11$ $3.87$$\pm 0.02$ $1.68$$\pm 0.01$ $2.07$$\pm 0.01$ $1.53$$\pm 0.02$
$K_{S}^{0}K^{-}\pi^{0}$ $(1.946,1.987)$ $(-40,31)$ $554$$\pm 50$ $7.51$$\pm 0.14$ $6.45$$\pm 0.09$ $8.73$$\pm 0.11$ $1.54$$\pm 0.02$ $0.68$$\pm 0.01$ $0.84$$\pm 0.01$ $0.53$$\pm 0.01$
$K^{-}\pi^{-}\pi^{+}$ $(1.953,1.983)$ $(-33,28)$ $1112$$\pm 83$ $23.47$$\pm 0.40$ $16.15$$\pm 0.17$ $21.19$$\pm 0.20$ $3.83$$\pm 0.03$ $1.66$$\pm 0.02$ $1.98$$\pm 0.02$ $1.46$$\pm 0.03$
$K_{S}^{0}K_{S}^{0}\pi^{-}$ $(1.951,1.986)$ $(-32,28)$ $266$$\pm 22$ $11.07$$\pm 0.22$ $7.96$$\pm 0.14$ $10.84$$\pm 0.18$ $1.94$$\pm 0.03$ $0.83$$\pm 0.02$ $1.03$$\pm 0.02$ $0.72$$\pm 0.02$
$K_{S}^{0}K^{+}\pi^{-}\pi^{-}$ $(1.953,1.983)$ $(-31,26)$ $808$$\pm 45$ $10.21$$\pm 0.12$ $7.46$$\pm 0.12$ $9.30$$\pm 0.14$ $1.80$$\pm 0.02$ $0.78$$\pm 0.02$ $0.88$$\pm 0.01$ $0.52$$\pm 0.02$
$K_{S}^{0}K^{-}\pi^{+}\pi^{-}$ $(1.958,1.980)$ $(-31,26)$ $390$$\pm 40$ $9.66$$\pm 0.20$ $6.70$$\pm 0.06$ $8.51$$\pm 0.07$ $1.64$$\pm 0.01$ $0.73$$\pm 0.01$ $0.79$$\pm 0.01$ $0.50$$\pm 0.01$
$\eta_{\gamma\gamma}\pi^{-}$ $(1.930,2.000)$ $(-52,43)$ $983$$\pm 69$ $19.33$$\pm 0.29$ $15.89$$\pm 0.08$ $22.99$$\pm 0.09$ $3.81$$\pm 0.02$ $1.63$$\pm 0.01$ $2.24$$\pm 0.01$ $1.69$$\pm 0.01$
$\eta_{\pi^{+}\pi^{-}\pi^{0}}\pi^{-}$ $(1.941,1.990)$ $(-43,34)$ $269$$\pm 29$ $11.22$$\pm 0.28$ $8.47$$\pm 0.06$ $11.92$$\pm 0.07$ $2.02$$\pm 0.01$ $0.90$$\pm 0.01$ $1.14$$\pm 0.01$ $0.79$$\pm 0.01$
$\eta\prime_{\pi^{+}\pi^{-}\eta}\pi^{-}$ $(1.940,1.996)$ $(-40,34)$ $575$$\pm 40$ $11.16$$\pm 0.18$ $8.60$$\pm 0.06$ $12.39$$\pm 0.07$ $2.05$$\pm 0.01$ $0.91$$\pm 0.01$ $1.18$$\pm 0.01$ $0.85$$\pm 0.01$
$\eta\prime_{\gamma\rho^{0}}\pi^{-}$ $(1.938,1.992)$ $(-43,33)$ $1233$$\pm 75$ $14.00$$\pm 0.19$ $11.04$$\pm 0.07$ $15.46$$\pm 0.08$ $2.68$$\pm 0.01$ $1.17$$\pm 0.01$ $1.49$$\pm 0.01$ $1.00$$\pm 0.01$
$\eta_{\gamma\gamma}\rho^{-}$ $(1.920,2.006)$ $(-66,49)$ $2142$$\pm 191$ $8.07$$\pm 0.13$ $7.64$$\pm 0.06$ $11.59$$\pm 0.07$ $1.83$$\pm 0.01$ $0.80$$\pm 0.01$ $1.13$$\pm 0.01$ $0.80$$\pm 0.01$
Table 2: The integrated luminosities, ${\mathcal{L}}$, the $M_{\rm BC}$ requirements, and the single-tag yields in data, $N_{\rm ST}$, for various energy points. The uncertainties are statistical only.
. $E_{\rm cm}$ (GeV) ${\mathcal{L}}$ (pb-1) $M_{\rm BC}$ (GeV/$c^{2}$) $N_{\rm ST}$ 4.237 530.3 $(2.107,2.117)$ $6477$$\pm 163$ 4.246 593.9 $(2.107,2.118)$ $11944$$\pm 246$ 4.260 828.4 $(2.107,2.118)$ $21550$$\pm 320$ 4.270 531.1 $(2.107,2.118)$ $13319$$\pm 244$ 4.280 175.7 $(2.106,2.119)$ $4063$$\pm 152$ 4.290 502.4 $(2.106,2.119)$ $9316$$\pm 221$ 4.310-4.315 546.3 $(2.106,2.119)$ $5758$$\pm 228$ 4.400 507.8 $(2.106,2.119)$ $1855$$\pm 87$ 4.420 1090.7 $(2.106,2.121)$ $14890$$\pm 443$ 4.440 569.9 $(2.106,2.121)$ $9699$$\pm 443$ 4.470-4.699 4768.3 $(2.104,2.123)$ $25156$$\pm 762$ Sum $124027$$\pm 1121$
V Analysis of leptonic $D^{+}_{s}$ decays
V.1 Selection of double-tag events
The candidates for $D^{*+}_{s}$ containing a leptonic $D^{+}_{s}$ decay are reconstructed from the unused showers and tracks remaining after the single-tag selection. We require only one additional track not used in the tag reconstruction: $N_{\rm extra}^{\rm charge}$= 0. The transition $\gamma(\pi^{0})$ is selected with the same selection criteria as the tag side.
Throughout this paper, the $\mu_{a}$ denote the $\mu^{+}$ candidates identified with MUC information, the $\mu_{b}$ denote the $\mu^{+}$ candidates identified without MUC information; and the $\tau^{+}_{e}$, $\tau^{+}_{\mu}$, $\tau^{+}_{\pi}$, and $\tau^{+}_{\rho}$ denote the $\tau^{+}$ candidates reconstructed via $\tau^{+}\to e^{+}\nu_{e}\bar{\nu}_{\tau}$, $\tau^{+}\to\mu^{+}\nu_{\mu}\bar{\nu}_{\tau}$, $\tau^{+}\to\pi^{+}\bar{\nu}_{\tau}$, and $\tau^{+}\to\rho^{+}\bar{\nu}_{\tau}$, respectively. In the selection of the candidates for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, the $\pi^{+}$ and $\rho^{+}$ candidates are selected with the same selection criteria as those used on the tag side. To select the candidates for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, the positron PID uses the measured information in the MDC, TOF, and EMC. The combined likelihoods ($\mathcal{L}^{\prime}$) under the positron, pion, and kaon hypotheses are obtained. Positron candidates are required to satisfy $\mathcal{L}^{\prime}(e)>0.001$ and $\mathcal{L}^{\prime}(e)/(\mathcal{L}^{\prime}(e)+\mathcal{L}^{\prime}(\pi)+\mathcal{L}^{\prime}(K))>0.8$. To select the candidates for $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, the muon candidate is required to have a deposited energy in the EMC within (0.1, 0.3)0.10.3(0.1,\,0.3)( 0.1 , 0.3 ) GeV. It must also satisfy a requirement on the hit depth, $d_{\mu^{+}}$, in the muon counter which depends on both $|\cos\theta_{\mu^{+}}|$ and $p_{\mu^{+}}$. These requirements on $d_{\mu^{+}}$ are shown in Table 3. The selected $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ candidates offer an opportunity to determine the BF of $D^{+}_{s}\to\mu^{+}_{b}\nu_{\mu}$ as a cross check.
Table 3: The $\cos\theta_{\mu^{+}}$ and $p_{\mu^{+}}$ dependent requirements on $d_{\mu^{+}}$ for muon candidates.
$|\cos\theta_{\mu^{+}}|$ $p_{\mu^{+}}$ (GeV/$c$) $d_{\mu^{+}}$ (cm)
$p_{\mu^{+}}\leq 0.88$ $d_{\mu^{+}}>17.0$
(0.00, 0.20) $1.040.88<p_{\mu^{+}}<$ 1.04 $d_{\mu^{+}}>100.0p_{\mu^{+}}-71.0$
$p_{\mu^{+}}\geq 1.04$ $d_{\mu^{+}}>33.0$
$p_{\mu^{+}}\leq 0.91$ $d_{\mu^{+}}>17.0$
(0.20, 0.40) $1.070.91<p_{\mu^{+}}<$ 1.07 $d_{\mu^{+}}>100.0p_{\mu^{+}}-74.0$
$p_{\mu^{+}}\geq 1.07$ $d_{\mu^{+}}>33.0$
$p_{\mu^{+}}\leq 0.94$ $d_{\mu^{+}}>17.0$
(0.40, 0.60) $1.100.94<p_{\mu^{+}}<$ 1.10 $d_{\mu^{+}}>100.0p_{\mu^{+}}-77.0$
$p_{\mu^{+}}\geq 1.10$ $d_{\mu^{+}}>33.0$
(0.60, 0.80) $d_{\mu^{+}}>17.0$
(0.80, 0.93) $d_{\mu^{+}}>17.0$
Information concerning the undetectable neutrino(s) is inferred by the kinematic quantity $M^{2}_{\mathrm{miss}}\equiv E^{2}_{\mathrm{miss}}/c^{4}-|\vec{p}_{\mathrm{miss}}|^{2}/c^{2}$, where $E_{\mathrm{miss}}$ and $\vec{p}_{\mathrm{miss}}$ are the missing energy and momentum of the neutrino(s) candidate, respectively, calculated by $E_{\mathrm{miss}}\equiv E_{\mathrm{beam}}-E_{k}$ and $\vec{p}_{\mathrm{miss}}\equiv-\vec{p}_{D_{s}^{*-}}-\vec{p}_{\gamma(\pi^{0})}-\vec{p}_{k}$ in the $e^{+}e^{-}$ center-of-mass frame. The index $k$ denotes the $e^{+}$, $\mu^{+}$, $\pi^{+}$, or $\rho^{+}$ of the signal candidate, and $E_{k}$ and $\vec{p}_{k}$ are its energy and momentum, respectively. For $D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}$, to improve the $M_{\rm miss}^{2}$ resolution, we perform a kinematic fit that constrains the masses of all possible particles and the missing neutrino combinations to the known mass of $D^{*-}_{s}$ and $D^{*+}_{s}$. If there is more than one combination, the one with the minimum $\chi^{2}$ values of the kinematic fit is retained for further analysis. The variable $E_{\rm sum}^{\rm extra\,\gamma}$ is defined as the total energy of good showers in the EMC including the transition $\gamma(\pi^{0})$ used to reconstruct the signal $D_{s}^{*+}$, but excluding those used in the tag side. We also remove bremsstrahlung photons candidates, defined as showers reconstructed within 10 degrees of the initial positron track direction.
Figure 2 shows the $M^{2}_{\rm miss}$ distribution of the candidate events for $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$. The left two plots of Fig. 3 show the $E_{\rm sum}^{\rm extra\,\gamma}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, and the right two plots of Fig. 3 show the $M^{2}_{\rm miss}$ distributions of the candidate events for $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. Different signal variables are chosen for the different $\tau^{+}$ modes mainly to achieve better separation between signal and background. These distributions are obtained from the combined data from all energy points.
V.2 Detection efficiencies
The detection efficiencies, $\epsilon_{\rm ST}^{ij}$ and $\epsilon_{\rm DT}^{ij}$, are estimated by analyzing MC samples with the relevant combinations of events and appropriate relative BFs. The single-tag MC sample is generated with inclusive $D^{*-}_{s}$ decays and $D^{-}_{s}$ decays to the single modes, while the double-tag has $D^{*-}_{s}$ decays to signal modes and $D^{-}_{s}$ decays to the tag modes. and $D^{*+}_{s}\to\rm anything$ with $D^{+}_{s}$ decays to signal modes. As an example, the double-tag efficiencies obtained for each signal decay at 4.260 GeV are shown in Table 1. For a given tag mode $i$ and energy point $j$, the effective signal efficiencies, $\epsilon_{\rm sig}^{ij}$, of each signal decay are obtained by dividing the $\epsilon_{\rm DT}^{ij}$ by $\epsilon_{\rm ST}^{ij}$. For each signal decay, the averaged signal efficiencies, $\bar{\epsilon}_{\rm sig}$, are obtained by weighting them by the relative single-tag yields $N_{\rm ST}^{ij}$; the results are shown in Table 4.
Figure 2: Fit to the $M_{\rm miss}^{2}$ distribution of the candidate events for $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ with the $\mu^{+}$ depth requirement. The points with error bars represent data combined from all energy points. The blue solid curve denotes the total fit. The red solid curves and red dashed curves show the fitted signal and combinatorial background shapes.
Figure 3: Simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions (left) of $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ as well as the $M_{\rm miss}^{2}$ distributions (right) of $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$. The points with error bars represent data combined from all energy points. The blue solid curve denotes the total fit. The red solid curves and red dashed curves show the fitted combinatorial background shapes. The pink dashed, cyan dashed and blue dashed are the backgrounds from $D_{s}^{+}\to\mu^{+}\nu_{\mu^{+}}$, $D_{s}^{+}\to K^{0}\pi^{+}$, $D_{s}^{+}\to\eta\pi^{+}\pi^{0}$ and $D_{s}^{+}\to K^{0}\pi^{+}\pi^{0}$, respectively.
V.3 Branching fractions without lepton universality constraint
When using the MUC information to identify muon candidates, the background level of $D_{s}^{+}\to\mu_{a}^{+}\nu_{\mu}$ is very low. In this case, the signal yield of $D_{s}^{+}\to\mu_{a}^{+}\nu_{\mu}$ is obtained from the unbinned maximum likelihood fit to the $M^{2}_{\rm miss}$ distribution, as shown in Fig. 2. In this fit, the signal and background shapes are modeled by the simulated shapes. We obtain a signal yield of $507\pm 26$. Combining this yield and detection, the efficiency, and the single-tag yield, we use Eq. 2 to obtain the BF of $D_{s}^{+}\to\mu_{a}^{+}\nu_{\mu}$ result:
$\mathcal{B}_{D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}}=(0.547\pm 0.026_{\rm stat}\pm 0.016_{\rm syst})\%.$
The systematic uncertainties in the BF measurement are discussed in the following section.
To extract the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, we perform a simultaneous fit to the $E_{\rm sum}^{\rm extra~{}\gamma}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{\ell}\tau_{\tau}$ ($\ell=e$ or $\mu$) and the $M_{\rm miss}^{2}$ distributions of candidates for $D^{+}_{s}\to\tau^{+}_{h}\tau_{\tau}$,( $h=\pi^{+}$ or $\rho^{+}$), as shown in Fig. 3. There is large component of $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$ in the accepted candidates $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ due to the poorer separation muons and pions when using only the $dE/dx$, TOF, and EMC information. The signal shapes of $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$, $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ are modeled by the individual simulated shapes for each decay mode. The signal shapes of the $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$ are described by a sum of two bifurcated-Gaussian functions, whose parameters are determined from the fits to the signal MC events and are fixed in the simultaneous fit. The background components for each signal decay are modeled with shapes derived from the relevant MC simulation The four $\tau^{+}$ decay modes are constrained to have a common BF for $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, taking into account different efficiencies and daughter particle decay BFs. From this simultaneous fit, we obtain
$\mathcal{B}_{D_{s}^{+}\to\mu^{+}_{b}\nu_{\mu}}=(0.493\pm 0.029_{\rm stat}\pm 0.020_{\rm syst})\%$
and
$\mathcal{B}_{D_{s}^{+}\to\tau^{+}\nu_{\tau}}=(5.60\pm 0.16_{\rm stat}\pm 0.20_{\rm syst})\%,$
which correspond to total signal yields of $579\pm 34$ and $2845\pm 83$, respectively. The systematic uncertainties in the BF measurements are discussed in the following section.
V.4 Branching fractions with lepton universality constraint
The ratio of decay widths between $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ and $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ is given by
$R=\frac{\Gamma_{D^{+}_{s}\to\tau^{+}\nu_{\tau}}}{\Gamma_{D^{+}_{s}\to\mu^{+}\nu_{\mu}}}=\frac{m^{2}_{\tau}(1-\frac{m^{2}_{\tau}}{m^{2}_{D_{s}}})^{2}}{m^{2}_{\mu}(1-\frac{m^{2}_{\mu}}{m^{2}_{D_{s}}})^{2}}.$ (5)
In this ratio, $f_{D^{+}_{s}}$ and $|V_{cs}|$ cancel, we can obtain a very precise SM prediction of $R=9.75\pm 0.01$, with the $\mu^{+}$ and $\tau^{+}$ masses from the PDG pdg2024 . To improve the precision of the measured $f_{D_{s}^{+}}|V_{cs}|$, we have examined the BFs of $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ after further constraining the ratio ${\mathcal{B}}(D^{+}_{s}\to\tau^{+}\nu_{\tau})/{\mathcal{B}}(D^{+}_{s}\to\mu^{+}\nu_{\mu})$ to be $9.75$ based on the SM prediction, in the fits to individual distributions in Fig. 3. From this constrained fit, the BFs of $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ and $D_{s}^{+}\to\tau^{+}\nu_{\tau}$ are obtained to be
$\mathcal{B}_{D_{s}^{+}\to\mu^{+}\nu_{\mu}}^{\rm SM}=(0.553\pm 0.014_{\rm stat}\pm 0.021_{\rm syst})\%,$
and
$\mathcal{B}_{D_{s}^{+}\to\tau^{+}\nu_{\tau}}^{\rm SM}=(5.39\pm 0.14_{\rm stat}\pm 0.20_{\rm syst})\%,$
respectively. These correspond to signal yields of $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ and $D_{s}^{+}\to\tau^{+}\nu_{\tau}$ to be $641\pm 16$ and $2754\pm 69$. The systematic uncertainties in the BF measurements are discussed in the following section.
Table 4: The signal yields, the effective signal efficiencies ($\bar{\epsilon}_{\rm sig}$), and the obtained BFs. The efficiencies include the BFs of all sub-resonant decays. The first uncertainties are statistical and the second systematic. The high effective signal efficiency for $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ is mainly due to that the tag environments in the signal MC sample is very different with the inclusive MC sample, where the single-tag efficiency is affected much due to low momentum photon(s) and pions.
Signal decay $\bar{\epsilon}_{\rm sig}$ (%) No lepton universality constraint Lepton universality constraint
[…]
$D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ $7.81\pm 0.02$ $2845\pm 83$ $5.60\pm 0.16\pm 0.20$ $2754\pm 69$ $5.39\pm 0.14\pm 0.20$
$D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ $18.57\pm 0.04$
$D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ $8.93\pm 0.02$
$D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$ $6.11\pm 0.02$
$D_{s}^{+}\to\mu^{+}_{b}\nu_{\mu}$ $94.76\pm 0.20$ $579\pm 34$ $0.491\pm 0.029\pm 0.020$ $641\pm 16$ $0.553\pm 0.014\pm 0.021$
$D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}$ $74.67\pm 0.16$ $507\pm 26$ $0.547\pm 0.026\pm 0.016$
VI Systematic uncertainties
The systematic uncertainties associated with the single-tag selection cancel. Several other sources of systematic uncertainties are estimated for the BFs measurements and described below.
VI.1 Individual systematic sources
VI.1.1 Single-tag yield
To estimate the systematic uncertainty in the fitted single-tag yield, we perform fits to the $M_{\rm BC}$ distributions of the data and inclusive MC sample with alternative signal and background shapes. The nominal signal shape is the simulated shape convolved with a Gaussian function. An alternative signal shape is obtained after requiring that the angle between each reconstructed and generated track is less than 20 degrees. The background shape is changed to a third-order Chebychev polynomial. The relative differences of the ST yields between data and the inclusive MC sample from these two variations are added in quadrature assigned as the systematic uncertainty. In addition, the uncertainty due to the background fluctuation of the fitted single-tag yield is considered as an additional systematic uncertainty. The quadrature sum of these two items, 1.88%, is assigned as the systematic uncertainty in the single-tag yield.
VI.1.2 Tracking and PID of $e^{+}$, $\mu^{+}$, and $\pi^{+}$
The efficiencies for tracking and PID of the $e^{+}$ and $\mu^{+}$ are studied with the control samples of $e^{+}e^{-}\to\gamma e^{+}e^{-}$ and $e^{+}e^{-}\to\gamma\mu^{+}\mu^{-}$, respectively. To consider the difference of topologies between $e^{+}e^{-}\to\gamma\ell^{+}\ell^{-}$ and $D^{+}_{s}\to\ell^{+}\nu_{\ell}$, the obtained efficiencies in different polar angles and momentum intervals of the control samples are re-weighted to match the $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ signals. The data-MC differences of the weighted efficiencies of $e^{+}$ tracking, and $e^{+}$ PID, $\mu^{+}$ tracking, and $\mu^{+}$ PID are summarized in Table 5. After correcting for the data-MC differences, the residual uncertainties are taken as individual systematic uncertainties, as listed in Table 6.
The efficiencies of the $\pi^{+}$ tracking and PID are studied with control samples of $e^{+}e^{-}\to K^{+}K^{-}\pi^{+}\pi^{-}(\pi^{0})$ and $\pi^{+}\pi^{-}\pi^{+}\pi^{-}(\pi^{0})$. The systematic uncertainties of the $\pi^{+}$ tracking and PID efficiencies are assigned to be 0.35% and 0.08%, respectively.
Table 5: Data-MC differences of the weighted efficiencies of tracking and PID for $e^{+}$ and $\mu^{+}$.
Signal decay $e^{+}$ tracking $e^{+}$ PID $\mu^{+}$ tracking $\mu^{+}$ PID
$D^{+}_{s}\to\mu^{+}_{a}\nu_{\mu}$ $99.93\pm 0.12$ $88.62\pm 0.63$
$D^{+}_{s}\to\mu^{+}_{b}\nu_{\mu}$ $99.93\pm 0.12$ $99.52\pm 0.53$
$D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$ $99.84\pm 0.08$ $98.45\pm 0.54$
$D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ $99.93\pm 0.12$ $89.05\pm 1.06$
VI.1.3 $\gamma(\pi^{0})$ selection
The photon selection efficiency was previously studied with $J/\psi\to\pi^{+}\pi^{-}\pi^{0}$ decays Upi0 . The $\pi^{0}$ reconstruction efficiency was previously studied with $e^{+}e^{-}\to K^{+}K^{-}\pi^{+}\pi^{-}\pi^{0}$ events. The systematic uncertainty of finding the transition $\gamma$ or $\pi^{0}$, weighted according to the BFs for $D_{s}^{*+}\to\gamma D_{s}^{+}$ and $D_{s}^{*+}\to\pi^{0}D_{s}^{+}$ pdg2024 , is 1.0%. For the $\pi^{0}$ in the $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, the systematic uncertainty is assigned to be 1.1% from the study of Ref. bes3_Ds_tauv1 . After re-weighting by the individual signal yields, the systematic uncertainty due to the $\pi^{0}$ in the $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$ for the overall BF measurement of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ is 0.29%.
VI.1.4 $E^{\mathrm{extra}~{}\gamma}_{\rm sum}$ and $N_{\rm extra}^{\rm charge}$ requirements
The efficiency for the requirements on $E^{\mathrm{extra}~{}\gamma}_{\rm sum}$ and $N_{\rm extra}^{\rm charge}$ is investigated with the double-tag sample of $D^{+}_{s}\to\eta\pi^{+}$. The ratio of the averaged efficiency of data to that of simulation is $1.058\pm 0.022$. After multiplying the signal efficiency by this factor, we assign 2.2% as the systematic uncertainty. Re-weighting by the individual signal yields, the systematic uncertainty due to $E^{\mathrm{extra}~{}\gamma}_{\rm sum}$ and $N_{\rm extra}^{\rm charge}$ requirements for the measurement of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ is 0.56%.
VI.1.5 Tag bias
The single-tag efficiencies obtained from the inclusive MC sample differ from those estimated with the signal MC events generated with events containing the single-tag $D_{s}^{*-}$ and signal decays, an effect known as “tag bias”. To estimate the MC simulation for these differences, we use a method from Ref. tagbias . To study the tag bias, we assign 1.0% for the tracking and PID efficiencies of $\pi^{+}$ and $K^{+}$, 2.0% for $\pi^{0}$, $K_{S}^{0}$, $\eta_{(\gamma\gamma)}$ reconstruction efficiencies of the tag side for their corresponding data and MC differences. The difference from 1 of $\varepsilon_{\rm ST}^{D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}}/\varepsilon_{\rm ST}^{\rm inclusive~{}D_{s}^{+}}$ is assumed to not cancel in the BF measurements. Weighting the offsets in each tag mode by their relative single-tag yield, the average offset for $D^{+}_{s}\to\mu_{a}^{+}\nu_{\mu}$ is calculated to be $(1.28\pm 0.01)\%$, which is taken as the systematic uncertainty.
Similarly, the systematic uncertainties due to tag bias are assigned as 1.19%, 1.28%, 1.22%, 1.19%, and 0.68% for $D^{+}_{s}\to\mu_{b}^{+}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, respectively. Re-weighting by individual signal yields, the systematic uncertainty due to tag bias for the measurement of the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ is $1.16\%$.
VI.1.6 MC statistics
The uncertainties due to the MC statistics are assigned to be 0.40%, 0.51%, 0.29%, 0.40%, and 0.32% for $D^{+}_{s}\to\mu^{+}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$, $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$, and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, respectively. Re-weighting by individual signal yields, the systematic uncertainty due to MC statistics for the measurement of the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ is $0.25\%$.
VI.1.7 $M^{2}_{\rm miss}$ and $E_{\rm sum}^{\rm extra~{}\gamma}$ fit
The uncertainties in the $M^{2}_{\rm miss}$ and $E_{\rm sum}^{\rm extra~{}\gamma}$ fits arise from the signal and background shapes.
For $D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}$, the systematic uncertainty due to the signal shape is estimated by replacing the nominal shape with signal shape convolved with a double Gaussian function with floating parameters. The relative change between the re-measured and nominal BFs, 1.12%, is taken as the systematic uncertainty. For $D^{+}_{s}\to\tau^{+}_{\mu}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{e}\nu_{\tau}$, we use the MC simulated shapes convolved with a single Gaussian resolution function with free parameters. For $D_{s}\to\mu_{b}\nu_{\mu}$, $D^{+}_{s}\to\tau^{+}_{\pi}\nu_{\tau}$ and $D^{+}_{s}\to\tau^{+}_{\rho}\nu_{\tau}$, the systematic uncertainty due to the signal shape is estimated by replacing the nominal shape by varying the Gaussian shape parameters individually by ±$\pm 1\sigma$. The quadrature sum of the relative changes between the re-measured BFs and the nominal BF, 0.60%, is assigned as the systematic uncertainty for $D^{+}_{s}\to\tau^{+}\nu_{\tau}$.
For $D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}$, the peaking background is mainly due to the misidentification of a $\pi^{+}$ or a $K^{+}$ as a $\mu^{+}$. We have corrected the background yields considering the data-MC differences. The systematic uncertainty is assigned by varying the weights of various background sources within ±$\pm 1\sigma$ of individual BFs. We also try alternative MC-simulated shapes by varying the relative fractions of the main backgrounds from $D_{s}^{\pm}D_{s}^{*\mp}$, $D_{s}^{*-}D_{s}^{*+}$ and $q\bar{q}$, by ±$\pm 1\sigma$ according to the observed cross sections Ref. crsDsDss and Ref. crsDssDss . After considering the uncertainties of the correction factors and changing the weights of different background shapes according to their yield uncertainties, we assign 0.80% as the associated systematic uncertainty.
To study the systematic uncertainty caused by the background shape in the fit for the measurement of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, we re-measure the BF by varying the background models in three ways. First, we use alternative MC-simulated shapes obtained by varying the relative fractions of the main backgrounds from $D_{s}^{\pm}D_{s}^{*\mp}$, $D_{s}^{*-}D_{s}^{*+}$ and $q\bar{q}$. We vary background yield by ±$\pm 1\sigma$ according to the observed cross sections crsDsDss and crsDssDss . Second, we try alternative MC-simulated shapes obtained by varying the relative fractions of their largest background sources. Third, we vary the yields of the main background sources by varying ±$\pm 1\sigma$ of the quoted BFs. Finally, we assign 2.34% as the systematic uncertainty in the measurement of the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$.
VI.1.8 Quoted branching fractions
The BFs of $\tau^{+}\to\mu^{+}\nu_{\mu}\bar{\nu}_{\tau}$, $\tau^{+}\to e^{+}\nu_{e}\bar{\nu}_{\tau}$, $\tau^{+}\to\pi^{+}\bar{\nu}_{\tau}$ and $\tau^{+}\to\rho^{+}\bar{\nu}_{\tau}$ are quoted from the PDG pdg2024 , which are $(17.39\pm 0.04)\%$, $(17.82\pm 0.04)\%$, $(10.82\pm 0.05)\%$ and $(25.49\pm 0.09)\%$, with relative uncertainties of 0.23%, 0.22%, 0.46%, and 0.35%, respectively. After re-weighting these by the individual signal yields, the systematic uncertainty for the BF of $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ is assigned as $0.14\%$.
The BFs of $D_{s}^{*+}\to\gamma D^{+}_{s}$ and $D_{s}^{*+}\to\pi^{0}D^{+}_{s}$ are ($93.5\pm 0.7$)and ($5.8\pm 0.7$)pdg2024 . By varying the BFs by these uncertainties, we find that the signal efficiencies change by no more than 0.30%, which is assigned as a systematic uncertainty. The effect of the BF uncertainty on $D_{s}^{*+}\to e^{+}e^{-}D^{+}_{s}$ is negligible.
The total systematic uncertainties for the quoted BFs of $D^{+}_{s}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ are 0.30% and 0.33%, respectively.
VI.2 Total systematic uncertainties without lepton universality constraint
Table 6 summarizes all systematic uncertainties in the measurements of the BFs of $D_{s}^{+}\to\mu_{a}^{+}\nu_{\mu}$, $D_{s}^{+}\to\mu_{b}^{+}\nu_{\mu}$, and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$. Assuming that all systematic uncertainties are independent with each other, the total systematic uncertainties in the measurements of the BFs of $D_{s}^{+}\to\mu_{a}^{+}\nu_{\mu}$, $D_{s}^{+}\to\mu_{b}^{+}\nu_{\mu}$, and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ are obtained by summing in quadrature, giving 2.88%, 4.13%, and 3.59%, respectively.
VI.3 Total systematic uncertainty with lepton universality constraint
When constraining the yields of $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$, the uncertainties of $N_{\rm ST}$ and quoted BFs are fully correlated and all other uncertainties are independent. For the independent system uncertainties, we vary the signal efficiency of the corresponding uncertainties ±$\pm 1\sigma$ to obtain the new BF measurement. The systematic uncertainty due to lepton universality constraint is assigned to be 0.04% by varying the fixed $R$ within ±$\pm 1\sigma$. We assign the relative change of the BF as the corresponding uncertainty. Table 6 summarizes the systematic uncertainties for individual sources. The total systematic uncertainty in the measurement of the BFs of $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ (SM) is the quadrature sum which gives 3.74%.
Table 6: Relative systematic uncertainties (%) in the measurements of the BFs of $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ and $D^{+}_{s}\to\tau^{+}\nu_{\tau}$.
Source $D_{s}^{+}\to\mu^{+}_{a}\nu_{\mu}$ $D_{s}^{+}\to\mu^{+}_{b}\nu_{\mu}$ $D^{+}_{s}\to\tau^{+}\nu_{\tau}$ $D^{+}_{s}\to\ell^{+}\nu_{\ell}$ (SM)
Single-tag yield 1.88 1.88 1.88 1.88
$\mu^{+}$ tracking 0.19 0.19 0.04 0.08
$\mu^{+}$ PID 0.63 1.06 0.21 0.99
$e^{+}$ tracking 0.11 0.10
$e^{+}$ PID 0.24 0.18
$\pi^{+}$ tracking 0.35 0.18
$\pi^{+}$ PID 0.08 0.04
$\gamma/\pi^{0}$ reconstruction 1.00 1.00 1.00 1.00
$\pi^{0}$ reconstruction from $\rho^{+}$ 0.29 0.19
$E^{\rm extra~{}\gamma}_{\rm sum}$ and $N^{\rm charge}_{\rm extra}$ requirements 0.40 0.56 0.37
$M^{2}_{\rm miss}$ fit and $E^{\rm extra~{}\gamma}_{\rm sum}$ fit 1.20 3.03 2.45 2.67
Tag bias 1.28 1.19 1.16 0.83
MC statistics 0.40 0.40 0.25 0.22
Quoted BFs 0.30 0.46 0.33 0.33
SM constraint 0.04
Total 2.88 4.13 3.59 3.74
VII CONCLUSION
By analyzing $10.64~{}\mathrm{fb}^{-1}$ of $e^{+}e^{-}$ collision data taken at $E_{\rm cm}$ between $4.237$ and $4.699$ GeV with the BESIII detector, the BFs of $D_{s}^{+}\to\mu^{+}\nu_{\mu}$ and $D_{s}^{+}\to\tau^{+}\nu_{\tau}$ are determined with and without constraining the ratios of their BFs to the SM prediction. The obtained BFs are shown in Table 4. Combining these BFs with the world averages of the masses of $\ell^{+}$ and $D_{s}^{+}$ as well as the lifetime of $D_{s}^{+}$, we obtain $f_{D_{s}^{+}}|V_{cs}|$ with Eq. 1. The comparison of the BFs and $f_{D_{s}^{+}}|V_{cs}|$ obtained in this work and the other measurements are shown in Table 7.
[…]
Table 7: Comparisons of the BFs and the corresponding products of $f_{D_{s}^{+}}|V_{cs}|$ from various experiments. The “weighted” values are obtained by combining results after considering the correlated effects. The “average” values are obtained by weighting both statistical and systematic uncertainties, but not the third uncertainty dominated by the uncertainty of the $D_{s}^{+}$ lifetime. The uncertainties of “average” BFs and the first uncertainties of “average” $f_{D^{+}_{s}}|V_{cs}|$ are the total experimental uncertainties combined from statistical and systematic effects, and the second uncertainties of “average” $f_{D^{+}_{s}}|V_{cs}|$ is due to the input uncertainty of the quoted lifetime of $D^{+}_{s}$. All “weighted” and “average” results do not include results constrained by the SM.
Experiment $E_{\rm cm}$ (GeV) Mode $D_{s}^{+}$ decay $\mathcal{B}$ (%) $f_{D_{s}^{+}}|V_{\rm cs}|$ (MeV)
This work 4.237-4.699 $\boldsymbol{D_{s}^{*\pm}D_{s}^{*\mp}}$ $\boldsymbol{\tau^{+}_{e}\nu_{\tau},\tau^{+}_{\mu}\nu_{\tau},\tau^{+}_{\pi}\nu_{\tau},\tau^{+}_{\rho}\nu_{\tau}}$ $\boldsymbol{5.60\pm 0.16\pm 0.20}$ $\boldsymbol{252.7\pm 3.6\pm 4.5\pm 0.6}$
This work (SM) 4.237-4.699 $\boldsymbol{D_{s}^{*\pm}D_{s}^{*\mp}}$ $\boldsymbol{\tau^{+}_{e}\nu_{\tau},\tau^{+}_{\mu}\nu_{\tau},\tau^{+}_{\pi}\nu_{\tau},\tau^{+}_{\rho}\nu_{\tau}}$ $\boldsymbol{5.39\pm 0.14\pm 0.20}$ $\boldsymbol{247.9\pm 3.2\pm 4.6\pm 0.5}$
BESIII bes4009 4.009 $D_{s}^{+}D_{s}^{-}$ $\tau^{+}_{\pi}\nu_{\tau}$ $3.28\pm 1.83\pm 0.37$ $193.4\pm 53.9\pm 10.9\pm 0.5$
BESIII bes3_Ds_tauv1 4.178-4.226 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\rho}\nu_{\tau}$ $5.30\pm 0.25\pm 0.20$ $245.8\pm 5.8\pm 4.6\pm 0.5$
BESIII bes3_Ds_tauv2 4.178-4.226 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\pi}\nu_{\tau}$ $5.21\pm 0.25\pm 0.17$ $243.7\pm 5.8\pm 4.0\pm 0.5$
BESIII bes3_Ds_tauv3 4.178-4.226 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{e}\nu_{\tau}$ $5.27\pm 0.10\pm 0.13$ $245.1\pm 2.3\pm 3.0\pm 0.5$
BESIII xiechenpiv 4.128-4.226 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\pi}\bar{\nu}_{\tau}$ $5.44\pm 0.17\pm 0.13$ $249.0\pm 3.9\pm 3.0\pm 0.5$
BESIII llcmvv 4.128-4.226 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\mu}\bar{\nu}_{\tau}$ $5.37\pm 0.17\pm 0.15$ $247.4\pm 3.9\pm 3.5\pm 0.5$
[…]
CLEO cleo2009b 4.170 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{e}\nu_{\tau}$ $5.30\pm 0.47\pm 0.22$ $245.8\pm 10.9\pm 5.1\pm 0.5$
CLEO cleo2009a 4.170 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\rho}\nu_{\tau}$ $5.52\pm 0.57\pm 0.21$ $250.9\pm 13.0\pm 4.8\pm 0.6$
CLEO cleo2009 4.170 $D^{\pm}_{s}D^{*\mp}_{s}$ $\tau^{+}_{\pi}\nu_{\tau}$ $6.42\pm 0.81\pm 0.18$ $270.5\pm 17.1\pm 3.8\pm 0.6$
BaBar babar2010 10.56 $DKX\gamma D^{-}_{s}$ $\tau^{+}_{e}\nu_{\tau},\tau^{+}_{\mu}\nu_{\tau}$ $4.96\pm 0.37\pm 0.57$ $237.8\pm 8.9\pm 13.7\pm 0.5$
Belle belle2013 10.56 $DKX\gamma D^{-}_{s}$ $\tau^{+}_{\pi}\nu_{\tau},\tau^{+}_{e}\nu_{\tau},\tau^{+}_{\mu}\nu_{\tau}$ $5.70\pm 0.21^{+0.31}_{-0.30}$ $254.9\pm 4.7\pm 7.0\pm 0.6$
Average $5.38\pm 0.09$ $247.7\pm 2.1\pm 0.5$
This work 4.237-4.699 $\boldsymbol{D_{s}^{*\pm}D_{s}^{*\mp}}$ $\boldsymbol{\mu^{+}_{a}\nu_{\mu}}$ $\boldsymbol{0.547\pm 0.026\pm 0.016}$ $\boldsymbol{246.5\pm 5.9\pm 3.6\pm 0.5}$
[…]
BESIII bes4009 4.009 $D_{s}^{+}D_{s}^{-}$ $\mu^{+}\nu_{\mu}$ $0.517\pm 0.075\pm 0.021$ $239.6\pm 17.4\pm 4.9\pm 0.5$
BESIII bes3_Ds_muv 4.178 $D_{s}^{\pm}D_{s}^{*\mp}$ $\mu^{+}\nu_{\mu}$ $0.549\pm 0.016\pm 0.015$ $246.9\pm 3.6\pm 3.4\pm 0.5$
BESIII bes3_Ds_tauv2 4.178-4.226 $D_{s}^{\pm}D_{s}^{*\mp}$ $\mu^{+}\nu_{\mu}$ $0.535\pm 0.013\pm 0.016$ $243.7\pm 3.0\pm 3.6\pm 0.5$
BESIII bes3_Ds_muv1 4.128-4.226 $D_{s}^{\pm}D_{s}^{*\mp}$ $\mu^{+}\nu_{\mu}$ $0.5294\pm 0.0108\pm 0.0085$ $242.5\pm 2.5\pm 1.9\pm 0.5$
[…]
CLEO cleo2009 4.170 $D^{\pm}_{s}D^{*\mp}_{s}$ $\mu^{+}\nu_{\mu}$ $0.565\pm 0.045\pm 0.017$ $250.5\pm 10.0\pm 3.8\pm 0.5$
BaBar babar2010 10.56 $DKX\gamma D^{-}_{s}$ $\mu^{+}\nu_{\mu}$ $0.602\pm 0.038\pm 0.034$ $258.6\pm 8.2\pm 7.3\pm 0.5$
Belle belle2013 10.56 $DKX\gamma D^{-}_{s}$ $\mu^{+}\nu_{\mu}$ $0.531\pm 0.028\pm 0.020$ $242.8\pm 6.4\pm 4.6\pm 0.5$
Average $0.539\pm 0.009$ $244.6\pm 2.0\pm 0.5$
VIII Acknowledgement
The BESIII Collaboration thanks the staff of BEPCII and the IHEP computing center for their strong support. This work is supported in part by National Key R&D Program of China under Contracts Nos. 2023YFA1606000, 2020YFA0406300, 2020YFA0406400, 2023YFA1606000; National Natural Science Foundation of China (NSFC) under Contracts Nos. 11635010, 11735014, 11935015, 11935016, 11935018, 11961141012, 12025502, 12035009, 12035013, 12061131003, 12192260, 12192261, 12192262, 12192263, 12192264, 12192265, 12221005, 12225509, 12235017; the Chinese Academy of Sciences (CAS) Large-Scale Scientific Facility Program; the CAS Center for Excellence in Particle Physics (CCEPP); Joint Large-Scale Scientific Facility Funds of the NSFC and CAS under Contract No. U1832207; 100 Talents Program of CAS; The Institute of Nuclear and Particle Physics (INPAC) and Shanghai Key Laboratory for Particle Physics and Cosmology; German Research Foundation DFG under Contracts Nos. 455635585, FOR5327, GRK 2149; Istituto Nazionale di Fisica Nucleare, Italy; Ministry of Development of Turkey under Contract No. DPT2006K-120470; National Research Foundation of Korea under Contract No. NRF-2022R1A2C1092335; National Science and Technology fund of Mongolia; National Science Research and Innovation Fund (NSRF) via the Program Management Unit for Human Resources & Institutional Development, Research and Innovation of Thailand under Contract No. B16F640076; Polish National Science Centre under Contract No. 2019/35/O/ST2/02907; The Swedish Research Council; U. S. Department of Energy under Contract No. DE-FG02-05ER41374.